log3 x-2 log3 2 = 3 Log3 3
help me solve this? p.s. the 3s' right next to the g's on log, are the base, sorry couldnt make them little threes.!

Answer by stanbon(75887)

log3 x-2 log3 2 = 3 Log3 3
-----
log3(x)-2log3(2) = 3log3(3)
-----
log3[x/2] = log3(27)
----
x/2 = 27
x = 54
============
Cheers,
Stan H.